Two point charges are placed as follows: charge q1 = -1.50 nC is at y = +6.00 m and charge q2 = +3.20 nC is at the origin. What is the total force (magnitude and direction) exerted by these two charges on a negative point charge q3 = -5.00 nC located at (2.00 m, -4.00 m)?

1
Expert's answer
2022-04-21T13:02:46-0400

"F=kq_3(\frac{q_1}{r_1^2}-\frac{q_2}{r_2^2})=2.5~nN."